OpenStudy (star_girl):

Write the equation: - the line with x-intercept 1 and y-intercept -2 in point slope form

OpenStudy (star_girl):

point-slope form \[y-y ^{1}=m (x-x ^{1})\]

Directrix (directrix):

Is the answer supposed to be in the form you posted?

OpenStudy (star_girl):

yup

Directrix (directrix):

IF the x intercept is 1, then the x intercept has coordinates (1,0) because to get on the x-axis, your y coordinate has to be 0. Similarly, if the y intercept is -2, then the y intercept has coordinates ( 0,-2).

Directrix (directrix):

So, I went over to the online equation of a line using two points calculator at http://www.webmath.com/equline1.html and cranked out the equation of the line. But, it is in slope-intercept form which is not a big deal. So, got over to http://www.webmath.com/equline1.html and enter the points. What follows is an explanation on how to get the equation of the line. But, look to see what the slope is cranked out to be.

Directrix (directrix):

That is a cool tool. For the point-slope form, choose one of the two points. Say, (0,-2) where x1 = 0 and y1 = -2 The online calculator cranked out the slope to be 2. Change in y divided by change in x. y - y1 = m * ( x - x1) y - (-2) = 2* (x - 0) --> point/slope form y + 2 = 2 * (x - 0) --> point/slope form simplified a bit.

OpenStudy (star_girl):

I got it thanks!! The answer is \[y-0=2(x-1)\.

OpenStudy (star_girl):

I'm sure I will use it later one. :)

Directrix (directrix):

I did not get that. Did you use the x-intercept.

OpenStudy (star_girl):

yup. Why??

Directrix (directrix):

My answer is the same as yours when both are placed in the y = mx + b form. In the point/slope form, you get a choice of which point to use which makes the answers look different when in reality they are the same.

OpenStudy (star_girl):

True.